Find the domain and range of f(x)=3/×^2+2
1
Expert's answer
2019-12-11T10:45:30-0500

"f(x)= \cfrac{3}{x^2}+2" .

Given that x is in the denominator, it cannot be equal to 0.

So, the domain of this function is "x \in (- \infty, 0)∪(0; + \infty)".

The value "\cfrac{3}{x^2}" could not be less than 0 because "x^2 >0" for all x.

The value  "\cfrac{3}{x^2}" could not be equal 0.

All of theese means that value "\cfrac{3}{x^2}+2 > 2" for all x.

So, the range of "f(x)" is "(2, + \infin )" .

Answer: Domain is "(- \infty, 0)∪(0, + \infty)" , range is "(2, + \infin )" .